# mmCooper: A Multi-agent Multi-stage Communication-efficient and Collaboration-robust Cooperative Perception Framework

### v1

@@ 3.4 Intermediate-stage Fusion
The data shared between agents inevitably contains calibration errors, which cause misalignment of feature maps. These errors may lead to the ego vehicle misjudging the positions of surrounding objects, thereby reducing detection performance. To address these issues, we implement a novel Multi-scale Offset-aware Fusion (MOF) Module in the intermediate-stage Module. By aggregating information from surrounding positions, the per-position feature fusion process introduces redundancy, allowing it to effectively tolerate calibration errors.
Multi-scale Offset-aware Fusion. The overall pipeline of this module is shown in Figure 4(a). First, the ego features and the received collaborator features are encoded into three scales to capture target information of varying sizes. Then, our multi-scale offset-aware attention is applied for feature fusion. Upsampling is performed on different scales to obtain fused features of the same size. The details of the Multi-scale Offset-aware Attention Module are shown in Figure 4(b). We encode the ego features into queries via a convolutional layer, and employ the spatially continuous features at the query positions of each collaborator’s features as keys and values in the cross-attention mechanism. Define the specified location as $(i,j)$ on the feature map, with a neighborhood size of $s\times s$. The output representation of feature fusion at location $c$ is expressed as:
$\mathcal{F}_{i,c}=CrossAttn(\sigma(F_{bev,c}^{i}),F_{bev,c}^{j,s\times s},F_{bev,c}^{j,s\times s}),$
[…]
@@ 3.5 Late-stage Fusion
Due to the limited perception capabilities of individual agents, they often generate imprecise or even incorrect bounding boxes. Existing late-stage fusion methods focus only on the fusion strategy at the current stage. By utilizing data sharing across multiple stages, we are able to fully utilize the abundant intermediate fused features to calibrate the received bounding boxes.
Figure 5: (a) The BBox Filtering & Calibration Module. (b) The Deformable Bounding Box Attention Module.
BBox Filtering & Calibration Module. In this module, we propose a Feature-guided Bounding Boxes Calibration Transformer (FBCFormer) to calibrate the bounding boxes transmitted by collaborators and filter out erroneous ones. As depicted in Figure 5(a), we use the received bounding boxes from collaborators and the intermediate-stage fused features as inputs to the Deformable Bounding Boxes Attention (DBA) Module. The remaining components follow the vanilla transformer setup [37]. We apply score and offset projections to the output of the FBCFormer to learn the scores and offsets for each bounding box, thereby removing incorrect bounding boxes and calibrating the coarse ones.
Figure 5(b) provides a detailed illustration of our DBA module. We first encode the bounding boxes received from collaborators. By applying our bounding boxes encoder to the bounding box coordinates, size, and orientation information, we effectively extract the bounding box feature. We use a deformable cross-attention module to focus on keypoints information, which yields a robust representation of bounding box errors. Specifically, the bounding box features serve as the initial query embedding $q$, and the bounding box coordinates act as reference points. A linear layer encodes each bounding box’s coordinates into position embeddings. Subsequently, the linear layer learns the 2D spatial offsets for each reference point, extracting keypoints information at $q+\bigtriangleup q$. The cross-attention layer aggregates keypoint information from multiple collaborator features to enhance the representation for each query embedding:
$DBA(q)=\sum_{\alpha=1}^{A}W_{\alpha}[\sum_{n=1}^{N}\sum_{m=1}^{M}\omega(W_{\beta}F_{b}(q))F_{j}(q+\bigtriangleup q_{m})],$
where $A$ is the number of attention heads, $W_{\alpha/\beta}$ are learnable parameters, $m$ represents the number of keypoints, and $\omega$ denotes the softmax operation. Finally, two linear layers are learned to determine the offsets and scores for each bounding box. The offsets are used to adjust the position, size, and angle of each bounding box, while the scores are filtered using a threshold selection function $f_{sel}(\cdot)$ to discard lower-scoring bounding boxes. To prevent learning unreasonable offsets and scores, we control the output range using the tanh activation function. Additionally, to train our FBCFormer component, we compute the offset loss $\mathcal{L}_{off}$ using the smooth absolute error loss between the detected bounding boxes and the corresponding ground truth bounding boxes with the same identifier. The score loss $\mathcal{L}_{score}$ is calculated using focal loss.
@@ 3.6 Loss Functions
We use two detection decoders to perform regression and classification on the fused features. The regression output represents the position, size, and yaw angle of each predefined box, expressed as $O_{reg}=f_{dec}^{r}(\mathcal{F}_{i})\in\mathbb{R}^{7\times H\times W}$. The classification output indicates the confidence score for each predefined box, represented as $O_{cls}=f_{dec}^{c}(\mathcal{F}_{i})\in\mathbb{R}^{2\times H\times W}$. To optimize the proposed system, we use smooth absolute error loss for regression and focal loss for classification, denoted as $\mathcal{L}_{reg}$ and $\mathcal{L}_{cls}$, respectively. Combining the above losses, our total loss is represented as:
$\mathcal{L}=\mathcal{L}_{reg}+\mathcal{L}_{cls}+\mathcal{L}_{off}+\mathcal{L}_{score}.$
Models Venue OPV2V DAIR-V2X V2XSet
Comm AP@ Comm AP@ Comm AP@
$(log2)(\downarrow)$ 0.7/0.5 ()($\uparrow$)( ↑ ) $(log2)(\downarrow)$ 0.7/0.5 ()($\uparrow$)( ↑ ) $(log2)(\downarrow)$ 0.7/0.5 ()($\uparrow$)( ↑ )
No Fusion* [15] CVPR 2019 - 48.66/68.71 - 43.57/50.03 - 40.20/60.60
Late Fusion* [15] CVPR 2019 10.36 59.48/79.62 9.08 34.47/51.14 10.26 30.75/54.92
Intermediate Fusion* [15] CVPR 2019 23.22 70.82/ 88.41 22.62 39.38/56.22 23.22 59.38/83.18
When2com [23] CVPR 2020 23.22 57.55/74.11 22.62 33.68/48.20 23.22 41.85/67.41
DisoNet [18] NIPS 2021 23.22 68.64/84.72 22.62 40.69/52.67 23.22 54.11/79.82
Where2comm [11] NIPS 2022 23.22 69.73/85.16 22.62 43.71/59.52 23.22 63.77/83.17
V2X-ViT [42] ECCV 2022 23.22 70.06/84.65 22.62 40.43/53.08 23.22 61.49/83.63
ERMVP [50] CVPR 2024 22.46 64.60/85.81 21.98 46.96/64.21 18.46 58.47/81.14
Ours - 10.09 78.11/ 88.93 15.33 48.27/ 65.12 9.06 65.86/84.40
Table 1: Collaborative perception performance on the OPV2V, DAIR-V2X and V2XSet dataset with a time delay of 100 $ms$, localization errors of 0.2 $m$, and heading errors of 0.2° using Average Percision(AP)@0.7/0.5 as metrics. $Comm$ denotes the communication volume. Bold numbers indicate the best results, while underlined numbers represent the second-best results. * represents a variation of PointPillar [15].
@@ 4.1 Datasets and experimental settings
Datasets. We conduct extensive validation of our proposed mmCooper on three benchmark datasets, namely OPV2V [43], DAIR-V2X [49] and V2XSet[42]. OPV2V is the first large-scale public V2V cooperative perception simulation dataset, collected using the autonomous driving simulator CARLA [8] and the cooperative driving simulation framework OpenCDA [41]. This dataset includes 6 types of road scenarios and 9 different cities. It comprises 73 different scenes, 11,464 frames of point clouds and RGB images, and over 230,000 annotated 3D detection boxes. The training, validation, and test sets are divided into 6,374, 1,980, and 2,170 frames, respectively. DAIR-V2X is a comprehensive large-scale real-world 3D object detection dataset designed for vehicle-road cooperative autonomous driving research. This dataset includes 71,254 frames of point cloud and image data, with the training, validation, and test sets split in a 5:2:3 ratio. The V2XSet dataset is a large-scale synthetic dataset designed for V2X perception. It consists of 73 scenarios collected using CARLA and OpenCDA. It contains a total of 11,447 frames, with each frame involving up to five agents. The dataset is divided into training, validation, and test sets, consisting of 6,694, 1,920, and 2,833 frames, respectively.
@@
Evaluation metrics. To evaluate the 3D object detection performance of the baseline and the proposed framework, we use the average precision (AP) [9] at Intersection-over-Union (IoU) thresholds of 0.5 and 0.7 as the evaluation metric. The communication volume between agents is expressed by the following equation:
$\textbf{V}_{comm}=\log_{2}{((\left|M\right|\times H\times W\times C+N_{bbx}\times 7)\times 32/8)},$
@@
where $\left|M\right|$ represents the feature retention ratio, $N_{bbx}$ denotes the number of bounding boxes per agent, and 7 is used to describe the coordinates, dimensions, and angle information of each bounding box. Each value is represented using the float32 data type, with division by 8 converting the result to bytes.
@@
Implementation details. We implement our model on the PyTorch toolbox [27] and train it on NVIDIA GeForce RTX 4090 GPUs. We use the Adam optimizer and adopt batch sizes of 3, 5, and 3 for the OPV2V, DAIR-V2X, and V2XSet datasets, respectively, with 60 epochs for each. All models use the PointPillars [15] backbone to extract features from the raw point clouds. The confidence ratio selected for the Confidence-based Filter Generation Module is 70% for OPV2V, 60% for DAIR-V2X, and 70% for V2XSet, and the surrounding grid size chosen for the Multi-scale Offset-aware Attention is 3×3. To closely simulate real traffic conditions, we introduce localization and heading errors with standard deviations of 0.2 $m$ and 0.2°, respectively, sampled from a Gaussian distribution into the initial state of each agent. The communication range is set to 70 $m$. Agents beyond the communication range do not participate in the collaborative process.
@@ 4.2 Quantitative Evaluation
Detection Performance Evaluation. Table 1 presents the comparison of the perception performance of our proposed mmCooper model and various baseline models on the OPV2V, DAIR-V2X and V2XSet datasets with a time delay of 100 $ms$, localization errors of 0.2 $m$, and heading errors of 0.2°. The baseline models included in the comparison are described as follows. The No Fusion model relies solely on local observations for 3D object detection. Late Fusion shares the perception results of individual predictions among agents. Intermediate Fusion allows agents to share complete features extracted from raw point clouds. The above three models are all variants based on the PointPillar[15] point cloud detector. Additionally, we consider state-of-the-art (SOTA) models, including When2com [23], DisoNet [18], Where2comm [11], V2XViT [42] and ERMVP [50]. As shown in the table, our proposed mmCooper outperforms previous methods on both simulated and real-world datasets, demonstrating the effectiveness of our approach. Compared to existing SOTA models, mmCooper outperforms the second-best baseline on the OPV2V, DAIR-V2X, and V2XSet datasets by 7.29%/0.59%, 1.31%/0.91% and 2.09%/0.77% in AP@0.7/0.5, respectively. Notably, mmCooper significantly reduces communication costs, achieving levels comparable to late fusion on both datasets. In comparison to other models, the communication volume required to achieve optimal performance is 9153/156/18305 times that of our method. On the OPV2V dataset, our method even achieves lower communication volume than late fusion. A possible explanation is that our Confidence-based Filter Generation Module first removes low-quality features and bounding boxes, then dynamically allocates intermediate-stage features and late-stage bounding boxes based on this. The dynamic allocation process prioritizes using bounding boxes to convey information, preserving a few key features while eliminating a large amount of background information.
@@
Table 2: Results for different localization errors ($m$) and transmission delay ($ms$) on the OPV2V and DAIR-V2X dataset evaluated using AP@0.7. Bold numbers indicate the best results, while underlined numbers represent the second-best results.
Robustness to Localization Errors and Transmission Delay. We verified the robustness of the agents during collaborative processes in the presence of localization errors and transmission delay on the OPV2V and DAIR-V2X datasets. Following the noise settings in [42], localization noise was sampled from a Gaussian distribution with a mean of 0 $m$ and a standard deviation $\sigma\in\{0,0.2,0.4\}m$. As shown in table 2, as the standard deviation of the localization noise gradually increases, the detection performance of the models correspondingly decreases. On the DAIR-V2X dataset, when2com even performs worse than the No Fusion method when the localization error exceeds 0.2 $m$. Our method outperforms previous SOTA models under all localization noise settings, effectively demonstrating its robustness to varying levels of localization error. Transmission delay can lead to misalignment of features and proposals, thereby impacting detection performance. We evaluated the models on the OPV2V and DAIR-V2X datasets with time delays $\tau\in\{0,200,400\}ms$ set to 0 $ms$, 200 $ms$, and 400 $ms$. With increasing time delay, the detection performance of all methods gradually degrades due to misleading features and proposals. However, mmCooper achieves higher Average Precision (AP) than current SOTA models on both datasets. This robustness to localization errors and transmission delay demonstrates that considering spatial continuity aligns features effectively, and intermediate-stage features provide significant guidance for subsequent bounding boxes.

### v3

@@ 3.4 Intermediate-stage Fusion
The data shared between agents inevitably contains calibration errors, causing potential misalignment in feature maps. To address this, we propose a Multi-scale Offset-aware Fusion (MOF) Module in the intermediate-stage fusion, which aggregates information from neighboring regions, allowing the fusion process at each location to account for potential misalignment. By incorporating the surrounding context, it effectively mitigates the impact of calibration errors, ensuring more accurate and robust agent collaboration.
Figure 5: (a) The BBox Filtering & Calibration (BFC) Module. (b) The Deformable BBox Attention (DBA) Module.
Multi-scale Offset-aware Fusion (MOF). The overall pipeline of this module is shown in Fig. 4(a). First, the ego features and received collaborator features are encoded at three different scales to capture target information of varying sizes. Next, Multi-scale Offset-aware Attention (MOA) is applied for feature fusion. The features are then upsampled to a uniform size and concatenated across scales to generate the final fused feature $F^{i}$.
Multi-scale Offset-aware Attention (MOA). As shown in Fig. 4(b), we encode the ego feature at each location into queries using MLP. Meanwhile, features from a small neighborhood around the corresponding positions in each collaborator’s feature map are extracted as keys and values for cross-attention. Note that low-confidence features are suppressed by collaborators and not transmitted, making them irrelevant during cross-attention. Specifically, we denote the focused location as $(r,c)$ on the feature map, with a neighborhood size of $s\times s$. The output representation of feature fusion at location $(r,c)$ is expressed as:
$\mathcal{F}_{r,c}^{i}=CrossAttn(MLP(F_{r,c}^{i}),\hat{F}_{r,c,s\times s}^{j},\hat{F}_{r,c,s\times s}^{j}),$
Where $\hat{F}_{r,c,s\times s}^{j}$ represents the transmitted collaborator features within an $s\times s$ neighborhood centered at position $(r,c)$. By selectively aggregating the features at this position and its neighborhood, we achieve collaboratively robust feature fusion.
@@ 3.5 Late-stage Fusion
The bounding boxes received from collaborators are often redundant, imprecise, misaligned, or even incorrect. To address this, the BBox Filtering & Calibration (BFC) Module in the Late-stage Fusion leverages the information-abundant fused intermediate features to refine the received bounding boxes, either filtering out erroneous ones or correcting their positions for improved accuracy.
BBox Filtering & Calibration (BFC). As depicted in Fig. 5(a), first, the received bounding boxes and the fused intermediate features will be put into the Deformable BBox Attention (DBA) Module as inputs. It then extracts relevant information for each received bounding box from collaborators, which is processed by a vanilla transformer [45] to generate quality scores and positional offsets. Note that, to prevent unreasonable offsets and scores, we bound the output range using the $Tanh$ activation function. Finally, bounding boxes with low quality scores are discarded, while the remaining ones are refined by applying the predicted positional offsets. The refined bounding boxes are then merged with those detected by the ego agent, followed by Non-Maximum Suppression (NMS) to generate the final results. To train the BFC module, we compute the offset loss $\mathcal{L}_{off}$ using the smooth absolute error loss [11] and the score loss $\mathcal{L}_{score}$ using focal loss [24].
Deformable BBox Attention (DBA). As shown in Fig. 5(b), we first encode the received bounding boxes $\hat{\mathcal{B}}^{j}$ by applying am MLP to their coordinates, size, and orientation information to generate feature representations. These features are then mapped onto the BEV space based on the bounding box locations, forming the bounding box feature map $F_{b}(q)$. To enhance these representations, we employ the deformable cross-attention mechanism to selectively attend to and integrate relevant information from the fused intermediate features. Specifically, the initial bounding box features $F_{b}(q)$ serve as the query embedding to calculate the offsets of reference points using MLP. Using these offsets, we retrieve the corresponding reference features from the fused intermediate feature map as $F^{i}_{r}=\mathcal{F}^{i}(q+\bigtriangleup q_{m})$. In the meanwhile, an MLP and Softmax are applied on $F_{b}(q)$ to generate the aggregation weights for $F^{i}_{r}$, which are then used to weight and integrate the reference features. Finally, after another MLP layer, we obtain the enhanced bounding box feature map:
$DBA(q)=\sum_{\alpha=1}^{A}W_{\alpha}[\sum_{n=1}^{N}\sum_{m=1}^{M}\omega(W_{\beta}F_{b}(q))F^{i}_{r}]+F_{b}(q),$
where $A$ is the number of attention heads, $W_{\alpha}$ and $W_{\beta}$ are learnable parameters, $m$ represents the number of reference locations, and $\omega$ denotes the softmax operation.
@@ 3.6 Loss Functions
We use the regression head and classification head of the Detection Decoder [4] to generate the bounding box detection results. The regression results represent the position, size, and yaw angle of each predefined box, expressed as $O_{reg}=f_{dec}^{r}(\mathcal{F}^{i})\in\mathbb{R}^{7\times H\times W}$. The classification results indicate the confidence score for each predefined box, represented as $O_{cls}=f_{dec}^{c}(\mathcal{F}^{i})\in\mathbb{R}^{2\times H\times W}$. To optimize the proposed system, we use smooth absolute error loss to supervise regression and focal loss for classification, denoted as $\mathcal{L}_{reg}$ and $\mathcal{L}_{cls}$, respectively. Combining the previous losses, our total loss is represented as:
$\mathcal{L}_{total}=\mathcal{L}_{reg}+\mathcal{L}_{cls}+\mathcal{L}_{off}+\mathcal{L}_{score}.$
@@ 4.1 Datasets and experimental settings
Datasets. We conduct an extensive evaluation on three benchmark datasets, namely OPV2V [50], DAIR-V2X [57], and V2XSet [49]. OPV2V is a large-scale public V2V cooperative perception simulation dataset, which comprises 73 different scenes, 11,464 frames of point clouds and RGB images, and over 230,000 annotated 3D detection boxes. The training, validation, and test sets are divided into 6,374, 1,980, and 2,170 frames, respectively. DAIR-V2X is a large-scale real-world 3D object detection dataset, including 71,254 frames of point cloud and image data, with the training, validation, and test sets split in a 5:2:3 ratio. The V2XSet dataset is a large-scale synthetic dataset designed for V2X perception. It contains a total of 11,447 frames, which is divided into training, validation, and test sets, consisting of 6,694, 1,920, and 2,833 frames, respectively.
@@
Evaluation metrics. To evaluate the 3D object detection performance of the baseline and the proposed framework, we use the average precision (AP) [10] at Intersection-over-Union (IoU) thresholds of 0.5 and 0.7 as the evaluation metric. The communication volume between agents is expressed by the following equation [13]:
$\textbf{V}=\log_{2}{((\left|M\right|\times H\times W\times C+N_{b}\times 7)\times 32/8)},$
@@
where $\left|M\right|$ represents the feature retention ratio, $N_{b}$ denotes the number of bounding boxes per agent, and 7 is used to describe the coordinates, dimensions, and angle information of each bounding box. Each value is represented using float32, with a division by 8 for bytes.
@@
Implementation details. We implement our model on the PyTorch toolbox [35] and train it on NVIDIA GeForce RTX 4090 GPUs. We use the Adam optimizer and adopt batch sizes of 3, 5, and 3 for the OPV2V, DAIR-V2X, and V2XSet datasets, respectively, with 60 epochs for each. All models use the PointPillars [18] backbone to extract features from raw point clouds. The confidence ratio $p$ selected for the CFG module is 70% for OPV2V, 60% for DAIR-V2X, and 70% for V2XSet, and the surrounding grid size chosen for the MOF is 3×3. To closely simulate real traffic conditions, we introduce localization and heading errors with standard deviations of 0.2 $m$ and 0.2°, respectively, sampled from a Gaussian distribution. The communication range is limited to 70mm, excluding agents beyond this range from collaborative communication.
Table 1: Collaborative perception performance on the OPV2V, DAIR-V2X and V2XSet dataset with a time delay of 100 $ms$, localization errors of 0.2 $m$, and heading errors of 0.2° using Average Percision(AP)@0.7/0.5 as metrics. Bold numbers indicate the best results, while underlined numbers represent the second-best results. * represents a variation of PointPillar [18].
Models OPV2V DAIR-V2X V2XSet
AP@0.7/0.5 AP@0.7/0.5 AP@0.7/0.5
No Fusion* [18] 48.66/68.71 43.57/50.03 40.20/60.60
Late Fusion* [18] 59.48/79.62 34.47/51.14 30.75/54.92
Intermediate Fusion* [18] 70.82/ 88.41 39.38/56.22 59.38/83.18
When2com [29] 57.55/74.11 33.68/48.20 41.85/67.41
DiscoNet [22] 68.64/84.72 40.69/52.67 54.11/79.82
Where2comm [13] 69.73/85.16 43.71/59.52 63.77/83.17
V2X-ViT [49] 70.06/84.65 40.43/53.08 61.49/83.63
Select2col [28] 62.46/82.30 34.82/51.96 51.22/76.88
ERMVP [59] 69.71/86.63 46.96/64.21 58.44/81.54
Ours 78.11/ 88.93 48.27/ 65.12 65.86/84.40
Figure 6: Collaborative perception performance and communication volumes of all the models on the OPV2V, DAIR-V2X, and V2XSet datasets.
@@ 4.2 Quantitative Evaluation
Detection Performance. Tab. 1 presents the comparison of the perception performance of our proposed mmCooper model and various baseline models on OPV2V, DAIR-V2X, and V2XSet datasets with a time delay of 100 $ms$, localization errors of 0.2 $m$, and heading errors of 0.2°. The No Fusion relies solely on ego agent observations. Late Fusion shares only the bounding box results among agents. Intermediate Fusion allows agents to share complete features extracted from raw point clouds. The above three models are all variants based on the PointPillar [18] point cloud detector. Additionally, we consider SOTA models, including When2com [29], DisoNet [22], Where2comm [13], V2X-ViT [49], Select2col [28] and ERMVP [59]. As shown in the table, our proposed mmCooper outperforms all the baselines on both simulated and real-world datasets. Compared to existing SOTA models, mmCooper outperforms the second-best baseline on the OPV2V, DAIR-V2X, and V2XSet datasets by 7.29%/0.59%, 1.31%/0.91% and 2.09%/0.77% in AP@0.7/0.5, respectively. This demonstrates the effectiveness and superiority of our multi-stage fusion framework.
Communication Volume. Fig. 6 illustrates the model performance under different bandwidth conditions on the OPV2V, DAIR-V2X, and V2XSet datasets. Experimental results show that mmCooper significantly reduces communication costs, achieving performance levels comparable to late fusion. In comparison to other SOTA models, the communication volume required to achieve optimal performance is 9153/156/18305 times more than our method. Notably, on the OPV2V and V2XSet datasets, our method even achieves lower communication volume than late fusion. This is because our CFG module suppresses low-confidence information from the transmission and selectively transmits complementary and exclusive information at intermediate and late stages rather than transmitting through only one stage encompassing all information of the entire scenario. In addition, since some Late Fusion methods (e.g., OpenCOOD [50]) transmit non-processed (e.g., NMS) BBoxes, our mmCooper even achieves a lower communication volume compared to the late fusion model on certain datasets.
Table 2: Results for different localization errors ($m$) and transmission delay (msms) on the OPV2V and DAIR-V2X dataset evaluated using AP@0.7. Bold numbers indicate the best results, while underlined numbers represent the second-best results.
@@
Robustness to Localization Errors and Transmission Delay. We verified the robustness of the agents during collaborative processes in the presence of localization errors and transmission delay on the OPV2V and DAIR-V2X datasets. Following the noise settings in [49], localization noise was sampled from a Gaussian distribution with a mean of 0 $m$ and a standard deviation $\sigma\in\{0,0.2,0.4\}m$. As shown in Tab. 2, as the standard deviation of the localization noise gradually increases, the detection performance of the models decreases. On the DAIR-V2X dataset, when2com even performs worse than the No Fusion method when the localization error exceeds 0.2 $m$. Our method outperforms all the baselines under all localization noise settings, demonstrating its robustness to varying levels of localization error. Transmission delay can lead to misalignment of features and proposals, thereby impacting detection performance. We evaluated the models on the OPV2V and DAIR-V2X datasets with time delays $\tau\in\{0,200,400\}ms$ set to 0 $ms$, 200 $ms$, and 400 $ms$. With increasing time delay, the detection performance of all methods gradually degrades. However, mmCooper achieves the highest Average Precision (AP) over all the models on both datasets. This robustness to localization errors and transmission delay demonstrates the effectiveness of the MOF and BFC modules in addressing the potential misalignment and errors during communication among agents. For more results regarding errors during transmission, please refer to the supplementary.